### v1

@@ 3.2 Long-form Video Understanding
We also evaluate our method on a comprehensive video understanding benchmark MVBench (Li et al., 2023b). In Table 3, our results could reveal that our mechanism will reserve the comprehensive video understanding ability over general video understanding tasks. Notably, our method with the same training data as PLLaVA, could achieve similar performance as 13B level model. We believe our method could obtain more information whether for short or long videos. To further validate the scalability of our model, we trained our method on the VideoChat2 (Li et al., 2023c) dataset. The results, illustrated at the bottom of Table 3, show that when trained on larger video datasets, VideoLLaMB improves accuracy on MVBench by 3.17 points and surpasses VideoChat2, which was trained on the same dataset.
@@ 3.3 Planning Tasks
Given the relatively brief duration of the input videos of the current planning benchmark, our comparative analysis includes both image-language and video-language models. The original protocol dictated the selection of a single frame corresponding to each action. To refine this approach and enhance the evaluation process, we introduce a smoother method. This involves segmenting the entire video into intervals based on predefined timesteps. This revised method is applied in the evaluation of the PLLaVA, LLaVA-NeXT-Video-DPO, and VideoLLaMB.
Results on EgoPlan (Chen et al., 2023)
The EgoPlan dataset (Chen et al., 2023) was developed as an egocentric question-answering benchmark tailored for embodied planning tasks, comprising 3,355 questions. The evaluation follows the framework established in the original study, utilizing the probability $p(a|v,l)$ to identify the most suitable answer candidates. In Table 4, we demonstrate that our model surpasses all other video-language models in performance. This suggests that our model’s use of memory significantly enhances its planning capabilities compared to methods focused on the current stage. While our approach does not outperform certain image-language models, we attribute this to the constraints of the current benchmark, which features brief action sequences and carefully curated frame-action pairs. Our goal is to develop more challenging benchmarks for egocentric planning and to adapt our method for real-time planning tasks. We are confident that our method holds great promise for generalizing to practical, real-world planning scenarios.
Qualatitive Results
@@
Image-Language Model
Qwen-VL-Chat (Bai et al., 2023) Qwen-7B 26.32
LLaVA-1.5 (Liu et al., 2023a) Vicuna-7B 26.80
SEED-LLaMA (Ge et al., 2023) LLaMA2-Chat-13B 29.93
InternLM-Xcomposer (Zhang et al., 2023c) InternLM-7B 34.4
Video-Language Model
VideoChatGPT (Maaz et al., 2023) LLaMA-7B 26.35
Valley (Luo et al., 2023) LLaMA-13B 26.17
VideoLLaMA (Zhang et al., 2023b) LLaMA2-Chat-7B 29.85
LLaVA-NeXT-Video (Zhang et al., 2024b) Vicuna-7B 28.96
PLLaVA (Xu et al., 2024b) Vicuna-7B 30.26
@@ 3.4 Streaming Caption
Streaming dense video captions (Chen et al., 2024b; Zhou et al., 2024) involves generating captions for videos in real-time, without the need to process the entire video sequence beforehand. The primary challenge in this task is determining the exact timestamps to predict event captions. Most existing methods rely on special tokens, annotated as the end of an action, for training. Our approach introduces the SceneTiling algorithm, which can automatically identify the break points in a streaming video and generate captions without requiring special training tokens. To enhance the efficiency of our method, we calculate the depth score using only the left similarity: $d_{i}=\left(cl(i)+s_{i}\right)/2$. In Figure 4, we present qualitative results of our method applied to a streaming video. These results demonstrate that our method can effectively detect scene changes and automatically generate event captions.
@@ 3.5 Stress Test: “Needle In a Video Haystack”
PLLaVA (Xu et al., 2024b) LLaVA-NeXT-Video-DPO (Zhang et al., 2024b)
Refer to caption Refer to caption
MA-LMM (He et al., 2024) LongVA (Zhang et al., 2024a)
@@
In NIAVH, we utilize ego-centric videos from the Ego4D (Grauman et al., 2022) dataset as the “haystack,” within which we seek to locate the “needle” provided in three distinct modalities: textual, image, and video. For the textual modality, we supply a crafted description; for the image modality, we use DALL-E [^4] https://openai.com/index/dall-e-3/ to create a corresponding image; and for the video modality, we employ Sora (Brooks et al., 2024) to generate a short video clip, all based on the same description. Each “needle” is set to a duration of 1 second and is inserted into the concatenated Ego4D videos at various depths and lengths. To evaluate the benchmark, a direct question about the details within the needles is set, and an LLM compares the response with the ground truth, providing a score from 1 to 10, with 10 indicating a perfect match. For quantitative results, we calculate the average scores for additional analysis.
Comparision with similar benchmarks  Recent work proposes a multimodal needle-in-a-haystack benchmark MM-NIAH (Wang et al., 2024a), which focuses on a mixture of images and documents as the haystack and only supports text and image needles. In contrast, NIVAH focuses on streaming video stacks and supports text, image, and video needles.
@@
In our experiment, we evaluate our approach with four distinct methods. These include (a) adaptive pooling (Xu et al., 2024b), (b) position extrapolation combined with sampling (Zhang et al., 2024b), and (c) the integration of resampler with memory retrieval and consolidation (He et al., 2024). (d) video alignment with long-context LLM without compression (Zhang et al., 2024a). For each model, we standardize the video frame rate to one frame per second, aligning the number of input frames with the duration of the video in seconds. This allows the inputs not to miss the needle information and all the models are in fair comparison. The outcomes of this evaluation are depicted in Figure 6. Our analysis leads to the following key observations:
  • Methods utilizing an adaptive pooling strategy risk omitting crucial information, as the length of the source material (the "haystack") is often many times greater than the target segment (the “needle”).
@@
  • Combining a resampler with a memory retrieval strategy markedly improves the encoding of extended information within a video. However, the length that can be encoded is ultimately constrained by the resampler’s compression capacity.
  • VideoLLaMB with memory retrieval is the most efficient at preserving previously encountered information. Nevertheless, it still exhibits shortcomings: it tends to forget earlier information and is prone to hallucination issues, such as misidentifying “holding book” as “holding phone”.
@@ 3.6 Performance Analysis
Our model’s recurrent strategy maintains a consistent visual input length to the LLM, significantly reducing GPU memory usage. While a larger memory cache theoretically requires more memory, the impact is minimal due to shorter memory tokens compared to visual input tokens. The recurrent memory operates on the bridge layer, minimizing intermediate costs. In our experiments on the EgoSchema (Mangalam et al., 2023) dataset, we compared our model against three video-language model categories: vanilla, pooling-based, and sampling-based. Results in Figure 7 show that our methods and other fixed-length input models significantly cut memory usage, with our approach compressing input length more effectively. Our design’s efficiency is evident, as the memory cache incurs negligible additional memory cost.
@@
did so even when compared to those employing a pooling strategy (Xu et al., 2024b). We attribute this improved performance to the efficient memory management mechanism integrated within the bridge layer of our method. This enables our approach to condense the visual input more effectively than others, resulting in shorter processing times of the LLM.
3.6.1 Ablation Study
@@
w.o. recurrent memory (mean pooling) 51.61 -2.19
w.o. recurrent memory (adaptive pooling) 49.4 -4.4
w.o. memory retrieval 52.2 -1.6
w.o. semantic segment (uniform segment) 52 -1.8
@@
In this section, we present an ablation study of our method, focusing on its individual components. We analysis our method the EgoSchema dataset. The corresponding results are detailed in Table 6. Initially, we assess the effectiveness of the recurrent mechanism. To do this, we replace this mechanism with two pooling strategies: mean pooling and adaptive pooling. For comparison purposes, we configure the adaptive pooling strategies to produce a target time sequence length of 4, matching our method’s settings. Our findings reveal that all pooling strategies cause a notable degradation in performance. Notably, the adaptive pooling strategy underperforms even mean pooling. We hypothesize that this discrepancy arises from differences in how training and inference are conducted; mean pooling, being more consistent, likely enhances the model’s generalizability. We then evaluate the memory retrieval mechanism and observe that it is indeed capable of preserving memory to a certain degree. Lastly, we examine the impact of our semantic segmentation strategy. Compared to a uniform segmentation approach, our method is more adept at dividing videos into semantic segments. This segmentation results in a more efficient preservation of information, mitigating the information loss typically associated with sampling strategies.
@@ 4 Related Work
The evolution of LLMs has significantly enhanced our ability to understand lengthy videos in terms of their interaction with human language. Methods for long video analysis fall into three categories: scaling-up approaches, agent-based techniques, and length extrapolation strategies. Scaling-up approaches focus on enlarging model parameters and extending training data (Liu et al., 2024a), or creating more efficient architectures to replace computationally intensive transformers (Li et al., 2024; Chen et al., 2024a), though these may not always be practical. Agent-based techniques utilize LLMs’ strategic planning, involving various visual experts for comprehensive understanding (Wang et al., 2023b; Choudhury et al., 2023; Fan et al., 2024) or converting visual inputs into textual descriptions (Wang et al., 2023c; Zhang et al., 2023a; Yang et al., 2024; Wang et al., 2024b), but can face efficiency issues or out-of-domain content challenges. Length extrapolation extends image-language and short video-language modeling to longer durations using strategies like temporal embeddings (Qian et al., 2024a), prompts (Ren et al., 2023), position encodings (Wang et al., 2024c; d), frame condensation (Song et al., 2023b), visual token compression (Korbar et al., 2023; Ma et al., 2023; Liu et al., 2024b), and retrieval-based methods with visual features (He et al., 2024), often through selective sampling which risks losing information. Our work introduces a recurrent memory strategy to encode entire video sequences and use a memory cache to preserve past memory, and project the memory-augmented current semantic segment into the LLM to maintain long video understanding ability.
@@
The field of planning, which entails the prediction of future actions based on past action sequences and the present context, has been substantiated as an effective approach within language models, as evidenced by several studies  (Driess et al., 2023; Song et al., 2023a; Mu et al., 2023). This methodology has parallels in video understanding, where the task of action anticipation based on visual data has gained traction  (Sener & Yao, 2019; Farha et al., 2020; Furnari et al., 2017). A burgeoning area of research is the intersection of action anticipation and goal-directed planning, which enhances the fundamental capabilities of artificial intelligence in the context of video understanding  (Patel et al., 2023; Zhao et al., 2023; Chen et al., 2023). This challenge is particularly acute in real-time streaming environments, where the system must not only interpret the current state but also retain a relatively extensive memory of past events to inform decision-making. Therefore, our proposed method could naturally suit this problem.
@@ 5 Conclusion
In conclusion, VideoLLaMB offers an advancement in video-language models by enhancing computational efficiency and efficacy. Utilizing Memory Bridge Layers with recurrent memory tokens and the SceneTilling algorithm, VideoLLaMB preserves crucial visual information and semantic coherence in long videos. The NIAVH benchmark robustly evaluates this capability. Empirical results show VideoLLaMB outperforms existing methods in long video QA, egocentric planning, and frame retrieval. In the future, we would like to integrate part of the LLM memory with the memory in the bridge while keeping the whole system efficient.
@@ Appendix A Parameter Analysis
We conducted a detailed parameter analysis of our model, focusing on two primary aspects: the number of memory tokens and the number of bridge layers. This analysis was performed using the EgoSchema dataset, under the experimental settings in Section B.1. The outcomes of this analysis are presented in Table 7. From the results, we observed a clear trend: a simultaneous increase in the number of memory tokens and the number of bridge layers leads to a notable improvement in performance. This finding is significant as it provides valuable direction for future enhancements of our method. To optimize our model further, we propose expanding the capacity of the bridge layer by adding more parameters while concurrently exploring more efficient architectural designs.
Appendix B Implementation Details
B.1 Implementation Details
In our experiment, we configured the memory tokens to a capacity of 32 and employed a single transformer layer as the bridge layer. For the training process, we set the number of training frames to 16 and limited the number of segments to 4. In order to ensure the visual encoder’s plug-and-play functionality, we froze its parameters, focusing the training solely on the bridge layer and the LLMs. We utilized the Image Encoder and Video Encoder from VideoLLaVA (Lin et al., 2023). In alignment with the procedures of PLLaVA (Xu et al., 2024b), we initialized the LLM using the LLaVA-1.5 (Liu et al., 2023a) configuration. The training dataset was identical to that used in PLLaVA, leveraging the same video data. To maintain the model’s proficiency in static visual learning, we retained the fine-tuning image data from LLaVA-1.5. Our experiments were conducted on four Nvidia A800 GPUs. Regarding other hyperparameters, we adhered to the original settings specified in the initialized models
B.2 Parameter Details
@@
B.3 Baseline Clarification
This work miss two long-video understanding model in some benchmarks for the following reasons: (1) the MALMM is built on InstructBLIP, which limits the input query length and, therefore, can’t be applied to the EgoSchma and the NExTQA benchmark. (2) MovieChat requires reloading the model at each test and requires heavy I/O pressure. Therefore, we only include the MALMM on our NIAVH benchmark for comparison.

### v2

@@ 3.3 Comprehensive Video Understanding
We evaluated our method using the comprehensive video understanding benchmark MVBench [32]. As shown in Table 3, our approach maintains strong performance across general video understanding tasks. Remarkably, with the same training data as PLLaVA, our method achieves performance comparable to a 13B-level model. Our method effectively extracts information from both short and long videos. To assess scalability, we trained our model on the VideoChat2 [30] dataset. The results, presented at the bottom of Table 3, demonstrate that our model, when trained on larger video datasets, improves accuracy on MVBench by 3.17 points and outperforms VideoChat2, which was trained on the same dataset.
@@ 3.4 Planning Tasks
The original protocol dictated the selection of a single frame corresponding to each action. To refine this approach and enhance the evaluation process, we introduce a smoother method. This involves segmenting the entire video into intervals based on predefined timesteps. This revised method is applied in the evaluation of PLLaVA, LLaVA-NeXT-Video-DPO, and VideoLLaMB.
Results on EgoPlan [9]
The EgoPlan dataset [9] was developed as an egocentric question-answering benchmark tailored for embodied planning tasks, comprising 3,355 questions. The evaluation follows the framework established in the original study, utilizing the probability $p(a|v,l)$ to identify the most suitable answer candidates. In Table 5, we demonstrate that our model surpasses all other video-language models in performance. This suggests that our model’s use of memory significantly enhances its planning capabilities compared to methods focused on the current stage. We are confident that our method holds great promise for generalizing to practical, real-world scenarios.
@@
VideoLLaMA [76] LLaMA2-Chat-7B 29.85
LLaVA-NeXT-Video [79] Vicuna-7B 28.96
PLLaVA [70] Vicuna-7B 30.26
@@ 3.5 Streaming Caption
Streaming dense video captions [7, 83] involves generating captions for videos in real-time, without the need to process the entire video sequence beforehand. The primary challenge in this task is determining the exact timestamps to predict event captions. Most existing methods rely on special tokens, annotated as the end of an action, for training. Our approach introduces the SceneTiling algorithm, which can automatically identify the break points in a streaming video and generate captions without requiring special training tokens. To enhance the efficiency of our method, we calculate the depth score using only the left similarity: $d_{i}=\left(cl_{i}-c_{i}\right)/2$. This demonstrates that our method can effectively detect scene changes and automatically generate event captions.
@@ 3.6 Stress Test: “Needle In a Video Haystack”
MA-LMM [20] LongVA [77]
@@
In NIAVH, we utilize ego-centric videos from the Ego4D [19] dataset as the “haystack,” within which we seek to locate the “needle” provided in three distinct modalities: textual, image, and video. For the textual modality, we supply a crafted description; for the image modality, we use DALL-E to create a corresponding image; and for the video modality, we use Sora [2] generated short video clip, all based on the same description. Each “needle” is set to a duration of 1 second and is inserted into the concatenated Ego4D videos at various depths and lengths. To evaluate the benchmark, a direct question about the details within the needles is set, and an LLM compares the response with the ground truth, providing a score from 1 to 10, with 10 indicating a perfect match. For quantitative results, we calculate the average scores for additional analysis.
Comparison with similar benchmarks  Recent work proposes a multimodal needle-in-a-haystack benchmark MM-NIAH [57], which focuses on a mixture of images and documents as the haystack and only supports text and image needles. In contrast, NIVAH focuses on streaming video stacks and supports text, image, and video needles.
@@
In our experiment, we evaluate our approach with four distinct methods. These include (a) adaptive pooling [70], (b) position extrapolation combined with sampling [79], (c) the integration of resampler with memory retrieval and consolidation [20], and (d) video alignment with long-context LLM without compression [77]. For each model, we standardize the video frame rate to one frame per second, aligning the number of input frames with the duration of the video in seconds. This ensures that the inputs contain the needle information and all the models are in fair comparison. The outcomes of this evaluation are depicted in Figure 3. Our analysis leads to the following key observations:
  • Methods utilizing an adaptive pooling strategy risk omitting crucial information, as the length of the source material (the ”haystack”) is often many times greater than the target segment (the “needle”).
@@
  • Combining a resampler with a retrieval strategy markedly improves the encoding of extended information in a video. However, the encoded length is ultimately constrained by the resampler’s compression capacity.
  • VideoLLaMB with retrieval is the most efficient at preserving previously encountered information. Nevertheless, it still exhibits shortcomings: it tends to forget earlier information and is prone to hallucination issues, such as misidentifying “holding book” as “holding phone”.
@@ 3.7 Performance Analysis
Our model’s recurrent strategy maintains a consistent visual input length to the LLM, significantly reducing GPU memory usage. While a larger memory cache theoretically requires more memory [66], the impact is minimal due to shorter memory tokens compared to visual input tokens. The recurrent memory operates on the bridge layer, minimizing intermediate costs. In our experiments on the EgoSchema [41] dataset, we compared our model against three categories: vanilla, pooling-based, and sampling-based. Results in Figure 4 show that our methods and other fixed-length input models significantly cut memory usage, with our approach compressing input length more effectively. Our design’s efficiency is evident, as the memory cache incurs negligible additional memory cost.
@@
did so even when compared to those employing a pooling strategy [70]. We attribute this improved performance to the efficient memory management mechanism integrated within the bridge layer of our method. This enables our approach to condense the visual input more effectively than others, resulting in shorter processing times of the LLM. We further analyze the composition of the inference time over videos of varying lengths, including the encoding time and generation time.
Video Duration Feature Process Generation All
30 0.25 1.65 1.9
300 2.3 1.91 4.21
3000 23.4 8.1 31.5
Table 7: Latency Analysis (sec.). We evaluate the inference time of different parts on different length videos.
3.8 Ablation Study
In this section, we present an ablation study of our method, focusing on its individual components. We analyze our method the EgoSchema dataset. The corresponding results are detailed in Table 8. Initially, we assess the effectiveness of the recurrent mechanism. To do this, we replace this mechanism with two pooling strategies: mean pooling and adaptive pooling. For comparison purposes, we configure the adaptive pooling strategies to produce a target time sequence length of 4, matching our method’s settings. Our findings reveal that all pooling strategies cause a notable degradation in performance. Notably, the adaptive pooling strategy underperforms even mean pooling. We hypothesize that this discrepancy arises from differences in how training and inference are conducted; mean pooling, being more consistent, likely enhances the model’s generalizability. We then evaluate the memory retrieval mechanism and observe that it is indeed capable of preserving memory to a certain degree. Lastly, we examine the impact of our semantic segmentation strategy. Compared to a uniform segmentation approach, our method is more adept at dividing videos into semantic segments. This segmentation results in a more efficient preservation of information, mitigating the information loss typically associated with sampling strategies.
@@ 3.8 Ablation Study
w.o. recurrent (mean pooling) 51.61 -2.19
w.o. recurrent (adaptive pooling) 49.4 -4.4
w.o. retrieval 52.2 -1.6
w.o. segment (uniform segment) 52 -1.8
@@ 4 Related Work
The advancement of large language models (LLMs) has significantly improved the comprehension of long videos through their interaction with human language. Current methodologies for long video analysis are categorized into scaling-up approaches, agent-based techniques, and length extrapolation strategies. Scaling-up approaches involve increasing model parameters and expanding training datasets [37], or developing more efficient architectures to replace computationally intensive transformers [31, 6], though these may not always be feasible. Agent-based techniques leverage LLMs’ strategic planning by incorporating various visual experts for comprehensive understanding [56, 10, 13, 63] or converting visual inputs into textual descriptions [59, 75, 72, 58], but they can encounter efficiency issues and challenges with out-of-domain content. Length extrapolation extends image-language and short video-language modeling to longer durations using techniques such as temporal embeddings [46], prompts [49], position encodings [64, 62], frame condensation [54], visual token compression [24, 39, 38], and retrieval-based methods with visual features [20]. These often involve selective sampling, which risks information loss. Our work introduces a recurrent memory strategy to encode entire video sequences, using a memory cache to retain past memory and project the memory-augmented current semantic segment into the LLM to maintain long video understanding.
@@
Anticipatory planning, which involves predicting future actions based on past sequences and current context, has been validated as effective in language models [12, 53, 42]. This approach is analogous to video understanding, where action anticipation based on visual data is gaining prominence [51, 14, 17]. A growing research area is the intersection of action anticipation and goal-directed planning, enhancing AI capabilities in video understanding [45, 80, 9]. This challenge is especially critical in real-time streaming environments, where systems must interpret the current state and retain an extensive memory of past events to inform decision-making. Our proposed method is well-suited to address these challenges.
@@ 5 Conclusion
VideoLLaMB presents a significant advancement in video-language modeling by improving both computational efficiency and long-context understanding. Through the introduction of Memory Bridge Layers with recurrent memory tokens and the SceneTiling algorithm, our approach effectively preserves essential visual cues and maintains semantic continuity across extended video sequences. Empirical evaluations show that VideoLLaMB consistently outperforms existing methods in tasks such as long VideoQA, egocentric planning, and frame retrieval. Looking ahead, we aim to explore the integration of LLM memory with the bridge memory, with a focus on preserving the system’s overall efficiency.
Acknowledgments  This work was supported by the grants from the National Natural Science Foundation of China (62376031, 62372014).
Supplementary Material
@@ Appendix A Parameter Analysis
We conducted a detailed parameter analysis of our model, focusing on two primary aspects: the number of memory tokens and the number of bridge layers. This analysis was performed using the EgoSchema dataset, under the experimental settings in Sec. C.1. The outcomes of this analysis are presented in Tab. 9. From the results, we observed a clear trend: a simultaneous increase in the number of memory tokens and the number of bridge layers leads to a notable improvement in performance. This finding is significant as it provides valuable direction for future enhancements of our method. To optimize our model further, we propose expanding the capacity of the bridge layer by adding more parameters while concurrently exploring more efficient architectural designs.
Appendix B Compression Strategy Analysis
In this section, we further explore the memory compression ability of our metho. We compare our method with two types trending memory compression methods: adaptive pooling  [70] and token compression  [54]. Different memory compression strategies are compared on the same LLM, training data, and compression rate for fair comparison. The results on Egoschema in Tab. 10 demonstrate our method could keep the memory in a better way.
Compression Strategy Accuracy
Adaptive Pooling [70] 45.6
Token Compression [54] 42.2
VideoLLaMB 53.8
Table 10: Comparison of different memory compression strategies.
Appendix C Implementation Details
C.1 Implementation Details
In our experiment, we configured the memory tokens to a capacity of 32 and employed a single transformer layer as the bridge layer. For the training process, we set the number of training frames to 16 and limited the number of segments to 4. In order to ensure the visual encoder’s plug-and-play functionality, we froze its parameters, focusing the training solely on the bridge layer and the LLMs. We utilized the Image Encoder and Video Encoder from Video-LLaVA [34]. In alignment with the procedures of PLLaVA [70], we initialized the LLM using the LLaVA-1.5 [36] configuration. The training dataset was identical to that used in PLLaVA, leveraging the same video data. To maintain the model’s proficiency in static visual learning, we retained the fine-tuning image data from LLaVA-1.5. Our experiments were conducted on four Nvidia A800 GPUs. Regarding other hyperparameters, we adhered to the original settings specified in the initialized models
C.2 Parameter Details
@@
C.3 Baseline Clarification
This work miss two long-video understanding model in some benchmarks for the following reasons: (1) the MALMM is built on InstructBLIP, which limits the input query length and, therefore, can’t be applied to the EgoSchma and the NExTQA benchmark. (2) MovieChat requires reloading the model at each test and requires heavy I/O pressure. Therefore, we only include the MALMM on our NIAVH benchmark for comparison. In addition, to make a fair comparison with different compression methods, we adopt these baselines in the same setting on Egoschema, and the results are illustrated on  Tab. 10.
Appendix D Qualitative Results
Planning
Figure 5: Qualitative results on EgoPlan.
Streaming Caption
In Figure Figure 6, we present the qualitative results of the streaming caption task. At the commencement of the video, the model is provided with the query: “Describe the video in one sentence”. Subsequently, at timestamps 0.00.00.0 seconds, 6.06.06.0 seconds, 8.08.08.0 seconds, and 10.010.010.0 seconds, the model autonomously generates captions in response to changes in the video scene, without requiring any user input.
D.1 Example of NIAVH
In this section, we visualize our proposed needle in a video haystack, which supports different modalities of needle, include text, image, and video. As is shown in Figure 7, the needle is “A young man is sitting on a piece of cloud in the sky, reading a book.”. For the text needle, we just append the text to the video directly; for the image and video needle, we insert the image and the video clips into the video haystack.